OpenStudy (anonymous):

Which of the following is the slope between the two points (1, 5) and (6, 4)? 1/5 -1/5 5 -5

OpenStudy (anonymous):

slope formula : m = y2-y1 over x2-x1 5-4 over 1-6

OpenStudy (dcolley):

 The formula is y2-y1/x2-x1

OpenStudy (anonymous):

thanks guys!

OpenStudy (anonymous):

\[\Huge y-y_1=\underbrace{\boxed{\frac{y_2-y_1}{x_2-x_1}}}_{m} (x-x_1)\] m=slope... lol

OpenStudy (anonymous):

anyone want to help me with some more?

OpenStudy (anonymous):

:)

OpenStudy (dcolley):

the same type? do you understand how to do it now?

OpenStudy (anonymous):

yes I understand now.. but no different types of Algebra 1 problems

OpenStudy (dcolley):

let's see

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

What is the slope and y-intercept of the graph of 3x + 2y = 12? a.slope -3/2 = ; y-intercept = (0, 6) b. slope = -3; y-intercept = (0, 12) c.slope -2/3 = ; y-intercept = (0, 6) d.slope = 6; y-intercept = (0, -3/2 )

OpenStudy (anonymous):

y intercept when x=0

OpenStudy (anonymous):

so eliminate options b and d right now....

OpenStudy (anonymous):

can you find slope now ? :) (without substituting x=0)...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now what

OpenStudy (anonymous):

have you solved it?

OpenStudy (anonymous):

is it A ?

OpenStudy (anonymous):

yes it is... :)

OpenStudy (anonymous):

Yay :)